# Understanding and Estimating the Execution Time of Quantum Programs

### v1

@@ 5.2 Model construction for execution time prediction
Figure 6: The neural network architecture of the model
@@
Fig. 6 shows the neural network architecture of the model we developed to predict the execution time of quantum circuits. For a quantum circuit, the dimension of its graph features is $n\times 178$, where $n$ represents the number of nodes, and each quantum circuit has a different number of nodes. Each set of graph features passes through three layers of Graph Transformer, and then through Global Average Pooling to reduce the dimensions to $1\times 178$. On the other hand, the dimension of a quantum circuit’s global features is $1\times 41$. The global features first pass through two fully connected layers, transforming their dimensions to $1\times 64$. These are then concatenated with the output from the Global Average Pooling. The dimensions of the concatenated results are $1\times 242$. This combined data is then fed into four additional fully connected layers to produce the final output, which is the predicted execution time of a quantum circuit on a backend, either a simulator or a quantum computer. When training the model, we utilized grid search to determine the optimal hyperparameters. As a result, we set the number of epochs to 500 and the batch size to 128.
5.3 Active learning for quantum execution time prediction
5.3.1 Active learning and greedy sampling techniques
@@
Active learning, a branch of machine learning, enables algorithms to selectively choose their training data, which is particularly useful when labels are costly or hard to obtain, in our case, collecting the execution time of quantum programs. This method enhances learning efficiency by maximizing accuracy with fewer labeled instances, thus reducing the overall cost of data labeling. It is especially relevant in scenarios where there is ample data but few labels [23]. Within active learning frameworks, Greedy Sampling (GS) plays a pivotal role in refining machine learning models.
@@
Greedy Sampling is a foundational technique where a model strategically chooses new samples to enhance accuracy rapidly with a minimal number of labeled instances, focusing on metrics such as informativeness, representativeness, and diversity to enrich the training dataset [36]. A brief summary of greedy sampling techniques is as follows:
@@
  • GS on the input domain ($G_{x}$): This approach increases diversity in the input space by first selecting samples nearest to the centroid of the unlabeled pool and then incrementally adding those furthest away, fostering a diverse training set [36].
@@
  • GS on the output domain ($G_{y}$): Concentrating on the output domain, $G_{y}$ aims to diversify the range of output values. It utilizes the predicted outputs from a regression model to gauge and incorporate diversity in the dataset [36].
@@
  • GS on input and output ($G_{xy}$): Integrating the strategies of $G_{x}$ and $G_{y}$, $G_{xy}$ augments diversity across both input and output spaces using a combined distance metric to optimize the selection of samples [36].
@@
5.3.2 Active learning for collecting execution time of quantum circuits
As described in Section 4, we selected 1510 quantum programs and two quantum computers (ibm_osaka and ibm_kyoto), constituting a dataset of 3020 ($1510\times 2$) data points. At a confidence level of 95% and an error margin of 5%, we need to sample 340 out of the 3020 available. For greedy sampling, we opt for the $G_{x}$ method, which iteratively selects samples based on the features of the input samples until the required sample size is achieved.
@@ 5.3.2 Active learning for collecting execution time of quantum circuits
$\begin{split}d_{x_{nm}}=\|x_{n}-x_{m}\|=\alpha\dot{\|}x_{n_{global}}-x_{m_{global}}\|+(1-\alpha)\dot{\|}x_{n_{graph}}-x_{m_{graph}}\|,\\ \quad m=1,\ldots,k;\quad n=k+1,\ldots,N;\quad\alpha\in[0,1]\end{split}$ (2)
@@
Algorithm 1 The GSx Active Learning method, modified from GS in [36, 37].
@@
5.4 Evaluation
5.4.1 Evaluation method
We begin by training the model on a dataset comprising execution times from 1510 quantum circuits across two simulators, yielding 3020 data points. This dataset is divided into training and test sets in a 9:1 ratio. We then select and save the best-performing model in the test set, characterized by the lowest MSE (Mean Squared Error). This model serves as a pre-trained model and is subsequently fine-tuned using data from real quantum computers. Next, we employ the saved pre-trained model to conduct 10-fold cross-validation on a dataset of 340 data points, chosen through greedy sampling from the dataset of execution times collected from quantum circuits executed on quantum computers. This cross-validation process involves randomly shuffling the dataset, splitting it into ten equal parts, designating one part as the test set, and using the remainder as the training set for each fold. We preserve the best result (i.e., the lowest MSE) from each fold. The average of these best results is then calculated and served as the final performance metric for predicting the execution times of quantum circuits on quantum computers.
5.4.2 Evaluation metrics
@@
$\text{NMSE}=\frac{\text{MSE}}{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (6)
@@
6 Results
This paper aims to understand and accurately predict the execution times of quantum programs on platforms such as the IBM Quantum Platform, through answering three research questions. In this section, we present our approaches and discuss the results for each research question.
6.1 RQ1: How do the execution times of quantum circuits vary across different circuits and backends?
6.1.1 Statistical analysis of quantum circuit execution times on simulators
@@
Table 1: Statistical information on quantum circuit execution times on simulators (in seconds)
@@
Figure 7: Histograms of execution time (in seconds) statistics for quantum circuits on different simulators
@@
The execution times of the same circuits on different simulators are statistically different but highly correlated. We perform statistical tests to compare and correlate the execution times of the circuits on the two simulators. Firstly, at a significance level of 0.05, we use the Shapiro-Wilk test [24] to determine if the execution times of quantum circuits on FakeSherbrooke and FakeWashington each follow a normal distribution. The Shapiro-Wilk test is a statistical method used to assess if a dataset is normally distributed. It is widely used to check the normality assumptions in data before applying certain statistical tests. The results of the Shapiro-Wilk Test are presented in the first two rows of Table 2, with $p$-values significantly less than 0.05. Consequently, it can be concluded that the execution times of quantum circuits are not normally distributed on both FakeSherbrooke and FakeWashington.
@@
Given the non-normal distributions, we then use the Wilcoxon Signed-Rank Test [15] to perform paired comparisons of the distributions of execution times of the same quantum circuit on the two different simulators. The Wilcoxon Signed-Rank Test is a nonparametric statistical test used to compare two related samples or matched pairs to determine whether the median difference between the pairs of observations is significantly different from zero. As shown in the third row of Table 2, the $p$-value is significantly less than 0.05. Hence, we conclude that there is a statistically significant difference between the execution times of the same quantum circuits on the two simulators.
@@
Finally, we employ Spearman’s Rank Correlation Coefficient measurement [26] to examine whether there is a correlation in the execution times of quantum circuits on the two different simulators. Specifically, we investigate whether the same quantum circuit tends to have either consistently higher or consistently lower execution times on both simulators. Spearman’s Rank Correlation Coefficient measures the strength and direction of a relationship between two variables using their ranks. It is useful when the data is not normally distributed or when dealing with ordinal variables. The result, a coefficient ranging from -1 to 1, shows whether the relationship is positive, negative, or non-existent. The results of the Spearman’s Rank Correlation Coefficient measurement are shown in the fourth row of Table 2. The $p$-value is significantly less than 0.05, leading us to conclude that there is a statistically significant correlation between the distributions. Moreover, with a $\rho$ value of 0.914, it is concluded that there is a strong positive correlation in the execution times of quantum circuits on the two simulators. Generally, this means that for the same quantum circuit, the execution times are consistently either both high or both low on the two simulators.
@@
Table 2: Statistical test results for comparing the circuits’ execution times on different simulators
@@
6.1.2 Statistical analysis of quantum circuit execution times on quantum computers
We follow a similar approach as in 6.1.1 to analyze the execution times of the circuits on real quantum computers. Table 3 and Fig. 8 present the distributions of the execution times of the quantum circuits on the two studied quantum computers.
@@ 6.1.2 Statistical analysis of quantum circuit execution times on quantum computers
Table 3: Statistical information on quantum circuit execution times (in seconds) on quantum computers
@@
Figure 8: Histograms of execution time (in seconds) statistics for quantum circuits on different quantum computers
@@
Table 4: Statistical test results for comparing the circuits’ execution times on different quantum computers
@@
6.1.3 Evaluation of the accuracy of execution times estimated by the IBM Quantum Platform
Given that quantum computing resources are limited and expensive, it underscores the necessity of accurately predicting the execution times of quantum circuits on quantum computers. The IBM Quantum Platform provides an estimation of the quantum execution time when a quantum circuit is submitted for execution. We first observe whether the estimtation provided by the IBM Quantum Platform meet our needs. When we submit a quantum circuit to a quantum computer provided by the IBM Quantum Platform for execution, we can obtain the predicted execution time of the quantum circuit by sending a GET request [^5] https://docs.quantum.ibm.com/api/runtime/tags/jobs. In this way, we recorded the estimated execution times of 192 quantum circuits executed at ibm_osaka and 148 at ibm_kyoto, provided by the IBM Quantum Platform. We then compare the estimated execution times with the actually measured execution time using the metrics described in 5.4.2. The results are presented in Table 5.
Table 5: Evaluation of IBM Quantum Platform’s execution time predictions

### v2

@@ 5.2. Model construction for execution time prediction
Figure 6. The neural network architecture of the model
@@
Fig. 6 shows the neural network architecture of the model we developed to predict the execution time of quantum circuits. For a quantum circuit, the dimension of its graph features is $n\times 178$, where $n$ represents the number of nodes, and each quantum circuit has a different number of nodes. Each set of graph features passes through three layers of Graph Transformer, and then through Global Average Pooling to reduce the dimensions to $1\times 178$. On the other hand, the dimension of a quantum circuit’s global features is $1\times 41$. The global features first pass through two fully connected layers, transforming their dimensions to $1\times 64$. These are then concatenated with the output from the Global Average Pooling. The dimensions of the concatenated results are $1\times 242$. This combined data is then fed into four additional fully connected layers to produce the final output, which is the predicted execution time of a quantum circuit on a backend, either a simulator or a quantum computer. When training the model, we utilized grid search to determine the optimal hyperparameters. Specifically, we split the dataset into training, validation, and test sets in an 8:1:1 ratio, with each partition kept fixed throughout the experiments. The training set is used to train the model, while the validation set is used to select the optimal hyperparameters. We vary the number of epochs from 100 to 800 in increments of 100, and the batch size from 32 to 160 in increments of 32. For each combination of hyperparameters, the model is retrained from scratch to avoid any data leakage. After evaluating performance across all combinations, we find that setting the number of epochs to 500 and the batch size to 128 produces the best results.
5.3. Active learning for quantum execution time prediction
5.3.1. Active learning and greedy sampling techniques
@@
Active learning, a branch of machine learning, enables algorithms to selectively choose their training data, which is particularly useful when labels are costly or hard to obtain, in our case, collecting the execution time of quantum circuits. This method enhances learning efficiency by maximizing accuracy with fewer labeled instances, thus reducing the overall cost of data labeling. It is especially relevant in scenarios where there is ample data but few labels (settles2009active). Within active learning frameworks, Greedy Sampling (GS) plays a pivotal role in refining machine learning models.
@@
Greedy Sampling is a foundational technique where a model strategically chooses new samples to enhance accuracy rapidly with a minimal number of labeled instances, focusing on metrics such as informativeness, representativeness, and diversity to enrich the training dataset (wu2019active). A brief summary of greedy sampling techniques is as follows:
@@
  • GS on the input domain ($G_{x}$): This approach increases diversity in the input space by first selecting samples nearest to the centroid of the unlabeled pool and then incrementally adding those furthest away, fostering a diverse training set (wu2019active).
@@
  • GS on the output domain ($G_{y}$): Concentrating on the output domain, $G_{y}$ aims to diversify the range of output values. It utilizes the predicted outputs from a regression model to gauge and incorporate diversity in the dataset (wu2019active).
@@
  • GS on input and output ($G_{xy}$): Integrating the strategies of $G_{x}$ and $G_{y}$, $G_{xy}$ augments diversity across both input and output spaces using a combined distance metric to optimize the selection of samples (wu2019active).
@@
5.3.2. Active learning for collecting execution time of quantum circuits
As described in Section 4, we selected 1510 quantum circuits and two quantum computers (ibm_osaka and ibm_kyoto), constituting a dataset of 3020 ($1510\times 2$) data points. At a confidence level of 95% and an error margin of 5%, we need to sample 340 out of the 3020 available. For greedy sampling, we opt for the $G_{x}$ method, which iteratively selects samples based on the features of the input samples until the required sample size is achieved.
@@ 5.3.2. Active learning for collecting execution time of quantum circuits
(3) $\begin{split}d_{x_{nm}}=\|x_{n}-x_{m}\|=\alpha\dot{\|}x_{n_{global}}-x_{m_{global}}\|+(1-\alpha)\dot{\|}x_{n_{graph}}-x_{m_{graph}}\|,\\ \quad m=1,\ldots,k;\quad n=k+1,\ldots,N;\quad\alpha\in[0,1]\end{split}$
@@
Algorithm 1 The GSx Active Learning method, modified from GS in (wu2019active; yun2019graph).
@@
5.4. Evaluation
5.4.1. Evaluation method
We begin by training the model on a dataset comprising execution times from 1510 quantum circuits across two simulators, yielding 3020 data points. This dataset is divided into training and test sets in a 9:1 ratio. We then select and save the best-performing model for the test set, characterized by the lowest MSE (Mean Squared Error). To minimize the use of costly real quantum hardware, this model serves as a pre-trained model and is subsequently fine-tuned using a small set of execution time data collected from real quantum computers. Next, we employ the saved pre-trained model to conduct 10-fold cross-validation on a dataset of 340 data points, chosen through greedy sampling from the dataset of execution times collected from quantum circuits executed on quantum computers. This process begins by randomly shuffling the dataset and dividing it into ten equal folds. For each fold, one part is designated as the test set, while the remaining nine folds (90% of the data) are further split into a training subset (80%) and a validation subset (10%) using a 8:1 ratio. We fix the hyperparameters to a batch size of 32 and 500 training epochs. After each training epoch, we evaluate the latest model on the validation subset (we obtain a new model after each epoch); finally, the best-performing model (i.e., the one with the lowest MSE on the validation subset) is selected. This selected model is then evaluated on the independent test set. The average of these results across the ten test folds is reported as the final performance metric for predicting the execution times of quantum circuits on quantum computers.
5.4.2. Evaluation metrics
@@
(7) $\text{NMSE}=\frac{\text{MSE}}{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$
@@
Note. While it is true that NMSE and $R^{2}$ are mathematically related by $R^{2}=1-\text{NMSE}$ under certain conditions, we intentionally include both metrics to improve readability and interpretability for different audiences. NMSE offers a normalized error scale that is directly comparable across datasets, while $R^{2}$ provides a goodness-of-fit measure commonly used in regression tasks. Reporting both helps highlight complementary perspectives on model performance.
6. Results
This paper aims to understand and accurately predict the execution times of quantum circuits on platforms such as the IBM Quantum Platform, through answering three research questions. In this section, we present our approaches and discuss the results for each research question.
6.1. RQ1: How do the execution times of quantum circuits vary across different circuits and backends?
6.1.1. Statistical analysis of quantum circuit execution times on simulators
@@
Table 1. Statistical information on quantum circuit execution times on simulators (in seconds)
@@
Figure 7. Histograms of execution time (in seconds) statistics for quantum circuits on different simulators
@@
The execution times of the same circuits on different simulators are statistically different but highly correlated. We perform statistical tests to compare and correlate the execution times of the circuits on the two simulators. Firstly, at a significance level of 0.05, we use the Shapiro-Wilk test (shapiro1965analysis) to determine if the execution times of quantum circuits on FakeSherbrooke and FakeWashington each follow a normal distribution. The Shapiro-Wilk test is a statistical method used to assess if a dataset is normally distributed. It is widely used to check the normality assumptions in data before applying certain statistical tests. The results of the Shapiro-Wilk Test are presented in the first two rows of Table 2, with $p$-values significantly less than 0.05. Consequently, it can be concluded that the execution times of quantum circuits are not normally distributed on both FakeSherbrooke and FakeWashington.
@@
Given the non-normal distributions, we then use the Wilcoxon Signed-Rank Test (Neuhäuser2011) to perform paired comparisons of the distributions of execution times of the same quantum circuit on the two different simulators. The Wilcoxon Signed-Rank Test is a nonparametric statistical test used to compare two related samples or matched pairs to determine whether the median difference between the pairs of observations is significantly different from zero. As shown in the third row of Table 2, the $p$-value is significantly less than 0.05. Hence, we conclude that there is a statistically significant difference between the execution times of the same quantum circuits on the two simulators.
@@
Finally, we employ Spearman’s Rank Correlation Coefficient measurement (spearman1987proof) to examine whether there is a correlation in the execution times of quantum circuits on the two different simulators. Specifically, we investigate whether the same quantum circuit tends to have either consistently higher or consistently lower execution times on both simulators. Spearman’s Rank Correlation Coefficient measures the strength and direction of a relationship between two variables using their ranks. It is useful when the data is not normally distributed or when dealing with ordinal variables. The result, a coefficient (i.e., the $\rho$ value) ranging from -1 to 1, shows whether the relationship is positive, negative, or non-existent. The results of the Spearman’s Rank Correlation Coefficient measurement are shown in the fourth row of Table 2. The $p$-value is significantly less than 0.05, leading us to conclude that there is a statistically significant correlation between the distributions. Moreover, with a $\rho$ value of 0.914, it is concluded that there is a strong positive correlation in the execution times of quantum circuits on the two simulators. Generally, this means that for the same quantum circuit, the execution times are consistently either both high or both low on the two simulators.
@@
Table 2. Statistical test results for comparing the circuits’ execution times on different simulators
@@
To investigate why the same quantum circuit yields different execution times on different simulators, we selected the circuit with the largest execution time discrepancy between the two simulators, namely qwalk-noancilla_indep_qiskit_9, for further analysis. As shown in Table 3, we present the compilation results of this circuit on both FakeSherbrooke and FakeWashington. The results reveal that the number of ECR (Echoed Cross-Resonance) gates in the FakeSherbrooke compilation is comparable to the number of Controlled-Not (CX) gates in the FakeWashington compilation. Since both ECR and CX are two-qubit gates, this part of the circuit complexity remains similar across the two backends. However, for other gate types—namely RZ, X, and SX—the circuit compiled for FakeSherbrooke contains a significantly higher number of gates compared to the compilation for FakeWashington. All gate-duration values are taken verbatim from each backend’s calibration snapshot [^6] https://quantum.cloud.ibm.com/docs/en/api/qiskit/0.44/qiskit.providers.models.BackendProperties. IBM realises an RZ gate as a virtual frame change, so its physical duration is effectively zero [^7] https://quantum.cloud.ibm.com/docs/en/api/qiskit/0.44/qiskit.circuit.library.RZGate. Moreover, single-qubit X-family gates (X, SX) share the same 160 dt DRAG pulse envelope, giving them identical calibrated durations (erata2024quantum). According to Table 4, both CX and ECR gates have similar average execution times, and their execution times are the longest among all gate types, which explains why the total execution times are high on both simulators. Nevertheless, due to the overwhelming number of SX and X gates introduced during the compilation for FakeSherbrooke, the overall execution time on FakeSherbrooke is significantly higher than that on FakeWashington.
Table 3. Compilation results of quantum circuit qwalk-noancilla_indep_qiskit_9 on two simulators
Metric FakeWashington FakeSherbrooke
Execution Time (s) 42.43 73.24
Gate Count: cx 18519
Gate Count: ecr 18273
Gate Count: rz 9099 48126
Gate Count: sx 125 36291
Gate Count: x 10 4262
Circuit Depth 19488 54509
Table 4. Average gate execution times on two simulators
Simulator cx (ns) ecr (ns) rz (ns) sx (ns) x (ns)
FakeWashington 550.41 0.0 35.56 35.56
FakeSherbrooke 533.83 0.0 56.89 56.89
We further quantify how target-independent features relate to execution time on simulators. Specifically, we employ Spearman’s rank correlation to examine, on the two simulators, whether execution time is associated with the target independent features, namely depth, num_qubits, two-qubit gate count, and num_gates. In addition, we use Pearson’s correlation coefficient (pearson1896vii) to assess whether these associations are linear. As shown in Table 5, depth is moderately associated both monotonically and linearly: FakeSherbrooke ($\rho=0.730$, $r=0.636$) and FakeWashington ($\rho=0.677$, $r=0.306$). Num_qubits exhibits only modest monotonic and negligible linear association: FakeSherbrooke ($\rho=0.382$, $r=0.081$) and FakeWashington ($\rho=0.352$, r=0.082r=-0.082). Two-qubit gate count shows a monotonic but limited linear association with time: on FakeSherbrooke (Spearman $\rho=0.796$, Pearson $r=0.315$) and on FakeWashington ($\rho=0.722$, $r=0.046$). Num_gates behaves similarly to two-qubit gate count: FakeSherbrooke ($\rho=0.795$, $r=0.355$) and FakeWashington ($\rho=0.722$, $r=0.062$). The weak linear relations of individual features indicate that single features are insufficient as predictors, so we need models that combine features. Moreover, the coefficients differ across backends; for example, on FakeWashington Pearson’s $r$ is generally smaller (e.g., 0.046 vs. 0.315 for the two-qubit gate feature) than on FakeSherbrooke, which suggests the need for backend dependent features and models.
Table 5. Monotonic vs. linear associations with execution time on simulators
Spearman $\rho$ Pearson $r$
Feature FakeSherbrooke FakeWashington FakeSherbrooke FakeWashington
Depth 0.730 0.677 0.636 0.306
Num_qubits 0.382 0.352 0.081 0.082-0.082
Two-qubit gate count 0.796 0.722 0.315 0.046
Num_gates 0.795 0.722 0.355 0.062
6.1.2. Statistical analysis of quantum circuit execution times on quantum computers
We follow a similar approach as in 6.1.1 to analyze the execution times of the circuits on real quantum computers. Table 6 and Fig. 8 present the distributions of the execution times of the quantum circuits on the two studied quantum computers. Careful readers may notice that the average execution time on the quantum computers is longer than that on the simulators. We would like to clarify that, in Figs. 7 and  8, for both simulators and real devices, ”execution time” denotes the backend-side system (quantum) time reported by IBM. It measures the time when the QPU is dedicated to a job and does not include queuing or host-side overhead [^8] https://quantum.cloud.ibm.com/docs/guides/max-execution-time. On real devices this duration includes hardware and control latencies, in particular qubit measurement and the per-shot repetition delay, which accumulate on the device side [^9] https://quantum.cloud.ibm.com/docs/guides/repetition-rate-execution; https://qiskit-community.github.io/qiskit-experiments/stable/0.6/manuals/measurement/restless_measurements.html. Simulators execute the same circuits in software and do not incur these hardware latencies, so in our workload their measured durations can be shorter.
@@ 6.1.2. Statistical analysis of quantum circuit execution times on quantum computers
Table 6. Statistical information on quantum circuit execution times (in seconds) on quantum computers
@@
Figure 8. Histograms of execution time (in seconds) statistics for quantum circuits on different quantum computers
@@
Table 7. Statistical test results for comparing the circuits’ execution times on different quantum computers
@@
Similarly, we examine how simple, target-independent features relate to execution time on quantum computers. On ibm_osaka and ibm_kyoto, Table 8 shows that depth is weak both monotonically and linearly: ibm_osaka ($\rho=0.246$, $r=0.220$) and ibm_kyoto ($\rho=0.221$, $r=0.231$). Num_qubits shows moderate association in both senses: ibm_osaka ($\rho=0.658$, $r=0.571$) and ibm_kyoto ($\rho=0.556$, $r=0.471$). Two-qubit gate count has weak-to-moderate monotonic and weak linear association with time: ibm_osaka ($\rho=0.355$, $r=0.291$) and ibm_kyoto ($\rho=0.275$, $r=0.212$). Num_gates shows weak to modest monotonic and linear associations: ibm_osaka ($\rho=0.201$, $r=0.242$) and ibm_kyoto ($\rho=0.309$, $r=0.369$). Taken together, the weak linear relations of individual features indicate that single features are insufficient as predictors, so we use learned models that combine features; in addition, the differences between backends suggest that backend dependent features and models are needed.
Table 8. Monotonic vs. linear associations with execution time on quantum computers.
Spearman $\rho$ Pearson $r$
Feature ibm_osaka ibm_kyoto ibm_osaka ibm_kyoto
Depth 0.246 0.221 0.220 0.231
Num_qubits 0.658 0.556 0.571 0.471
Two-qubit gate count 0.355 0.275 0.291 0.212
Num_gates 0.201 0.309 0.242 0.369
6.1.3. Evaluation of the accuracy of execution times estimated by the IBM Quantum Platform
Given that quantum computing resources are limited and expensive, it underscores the necessity of accurately predicting the execution times of quantum circuits on quantum computers. The IBM Quantum Platform provides an estimation of the quantum execution time when a quantum circuit is submitted for execution. It is important to note that the execution‑time estimates provided by the IBM Quantum Platform represent worst‑case values for the circuit execution itself and do not include any queueing delays. We first observe whether the estimtation provided by the IBM Quantum Platform meets our needs. When we submit a quantum circuit to a quantum computer provided by the IBM Quantum Platform for execution, we can obtain the predicted execution time of the quantum circuit by sending a GET request [^10] https://docs.quantum.ibm.com/api/runtime/tags/jobs. In this way, we recorded the estimated execution times of 192 quantum circuits executed at ibm_osaka and 148 at ibm_kyoto, provided by the IBM Quantum Platform. We then compare the estimated execution times with the actually measured execution time using the metrics described in 5.4.2. The results are presented in Table 9.
Table 9. Evaluation of IBM Quantum Platform’s execution time predictions